# Debugging Concept Bottleneck Models through Removal and Retraining

### v1

@@ 1 Introduction
Concept Bottleneck Models (CBMs) (Koh et al., 2020) have emerged as a powerful architecture for interpretable vision classification. A CBM consists of two stages: a concept extractor first predicts a set of human-understandable concepts, which are then passed to an inference layer to produce the final label. This intermediate representation allows domain experts to inspect the model’s reasoning process and verify whether it aligns with their own. This capability is crucial in high-stakes domains, such as healthcare or scientific analysis, where errors are costly and expert validation is essential.
Beyond passive validation, CBMs enable test-time interventions (Koh et al., 2020; Espinosa Zarlenga et al., 2023). An expert can review the predicted concepts and directly correct them to influence the final prediction. For example, if a radiologist corrects a mispredicted concept, such as marking a lesion as present when the model missed it, the diagnosis may shift from benign to malignant. Such interventions elevate the expert from a passive auditor to an active participant in decision making. However, the effectiveness of these interventions hinges on a critical assumption: the learned concepts must align with expert knowledge. In practice, this alignment is often fragile. Data quality issues, sampling bias, and incomplete concept vocabularies can lead models to exploit spurious correlations or overlook key factors (Torralba & Efros, 2011; Geirhos et al., 2020). Because test-time edits address only surface errors, the same reasoning flaws inevitably reappear on new samples.
@@
In this work, we present a general debugging framework for CBMs that enables experts to globally edit a model’s reliance on undesired concepts, ensuring that its predictions are not only accurate but also right for the right reasons Ross et al. (2017), and aligned with the domain expert’s reasoning. This framework follows a two-step process of Removal and Retraining (Figure 1).
@@ 2 Related Work
These models fall into two main architectural families. ProtoPNets (Chen et al., 2019) and their extensions (Nauta et al., 2023b; Ma et al., 2024; Carmichael et al., 2024), learn prototypical patches from the training set to make predictions. In this family, concept-level explanations take the form of representative image patches from the training data. VLM-CBMs (Debole et al., 2025) use a CLIP backbone to score the presence of textual concepts, as explored in many approaches (Yuksekgonul et al., 2023; Oikarinen et al., 2023; Srivastava et al., 2024), where concept-level explanations are the textual descriptions of each concept. Our framework can also extend to post-hoc XAI methods (e.g., interpreting neurons from class activation maps Zhou et al. (2016) as concepts). In this work we focus on concept bottlenecks, which are interpretable-by-design, and evaluate one representative from each family: PIP-Net (Nauta et al., 2023b) and Post-hoc CBM (Yuksekgonul et al., 2023).
Interpretable Debugging. The goal of interpretable debugging is to enable a domain expert to interact with an interpretable model to detect and correct undesired behaviors. Early work focused on explanatory interactive learning (Ross et al., 2017; Teso & Kersting, 2019; Schramowski et al., 2020), with subsequent extensions to neuro-symbolic models (Stammer et al., 2022; 2021). Bontempelli et al. (2023) advanced these approaches, focusing on ProtoPNets and removing the need for a fixed concept vocabulary. Building on top of these works, we focus on a generalized unsupervised CBM architecture (Poeta et al., 2023).
Within unsupervised CBMs, spurious concept removal has been explored in both ProtoPNets and VLM-CBMs: Nauta et al. (2023a) study removal in PIP-Net on medical datasets, while Yuksekgonul et al. (2023) and Rao et al. (2024) evaluate similar strategies for VLM-CBMs. Retraining efforts have focused mainly on ProtoPNets, such as adding object segmentation maps for extra supervision (Barnett et al., 2021) or ProtoPDebug (Bontempelli et al., 2023) which applies a forgetting loss to marked concepts. Donnelly et al. (2025) bypasses the need for retraining, but is limited to random patch selection for learning new concepts. For VLM-CBMs, Bontempelli et al. (2021) outline debugging strategies but lack empirical evaluation. We build on these directions with a general debugging framework for CBMs and an effective retraining approach, CBDebug, that leverages a novel connection between interpretable debugging and bias mitigation to remove undesired concepts and better align models with expert feedback.
@@
Supervised methods require auxiliary labels to reduce the impact of spurious correlations (Zheng & Makar, 2022; Sagawa et al., 2019; Makar et al., 2022; Kirichenko et al., 2022). While these approaches cannot be directly applied to unsupervised CBMs, CBDebug utilizes a crucial connection between interpretable debugging and these methods. By leveraging the model’s interpretability, we can bridge concept-level human feedback to sample-level auxiliary labels, effectively removing the reliance on any concept marked as undesired by the domain expert. This is accomplished by collecting the activation scores for each concept on every sample, which can then be used to form the auxiliary labels. Specifically, we instantiate our approach with permutation weighting (Arbour et al., 2021), which can handle high-dimensional, real-valued auxiliary labels. This allows us to use concept activations directly, unlike methods such as GroupDRO (Sagawa et al., 2019) that require discrete groups and would necessitate an additional clustering step to convert activation scores.
@@ 3.2 Removal
To guide concept removal, we adopt a simple binary feedback mechanism: each concept is either retained or marked for removal, based on expert input. This minimal supervision design ensures that our approach remains broadly applicable and easy to integrate into real-world workflows, where experts may have limited time or domain knowledge to provide detailed annotations.
@@ 4.2 Reweight
We adopt permutation weighting (Arbour et al., 2021), later applied to shortcut removal by Zheng & Makar (2022), to perform reweighting. Unlike group-based reweighting approaches such as GroupDRO (Sagawa et al., 2019), which require discrete group labels and often struggle with groups that have low support (necessitating clustering of $\smash{\mathbf{\widehat{V}}}$), permutation weighting naturally accommodates multi-dimensional, continuous auxiliary labels. By directly enforcing independence between $\smash{\mathbf{\widehat{V}}}$ and $Y$, it provides a more general and stable mechanism.
@@
where $\eta(y_{i},v_{i})$ denotes the estimated probability that a sample belongs to $D^{\prime}$. To ensure robust weights, we perform K-fold cross validation and average over multiple permutations.
@@ 4.3 Augment
Samples assigned a low weight are more likely to be aligned with the bias we want to remove, and so we would like to focus our augmentation on these samples. To accomplish this, we can convert each sample weight $u_{i}$ into an augmentation probability $p_{aug}(x_{i})$. We first invert the sample weight $u_{i}$ by subtracting each weight from the maximum sample weight assigned. Then we normalize the resulting values to $[0,1]$ to convert them into probabilities, and raise the probabilities to a power $\gamma$ to increase contrast and reduce the likelihood of augmenting useful samples.
We then augment each sample with probability $p_{aug}(x_{i})$. Our augmentation strategy is dependent on the concept representation: For ProtoPNets we randomly select $k$ spurious concepts from $\mathcal{C}_{spur}$ and perform CutMix (Yun et al., 2019) using one of the top ten most activated patches for each concept. For VLM-CBMs we randomly select a spurious concept from $\mathcal{C}_{spur}$ and perform Mixup (Zhang, 2017) with an image of that concept selected from a text-to-image–generated concept bank, following DISC (Wu et al., 2023).
@@ 5 Experiments
Datasets. We use datasets with known spurious correlations: Waterbirds (Sagawa et al., 2019), MetaShift (Liang & Zou, 2022), CelebA (Liu et al., 2015), and ISIC (Codella et al., 2019). These datasets provide concrete testbeds for assessing how well CBDebug reduces reliance on undesired concepts, as their group structures allow performance to be measured directly across subpopulations.
@@ 5.1 Can CBDebug effectively retrain based on user feedback?
CBDebug. In contrast to these naive baselines, CBDebug improves worst-group accuracy by 7.5% on Waterbirds and 4.9% on MetaShift for PIP-Net, and by 26.1% on Waterbirds and 4.8% on MetaShift for Post-hoc CBM. CBDebug surpasses the previous state-of-the-art interpretable debugger, ProtoPDebug, while integrating its component steps into a framework that delivers more stable gains across settings. These consistent improvements across architectures and datasets highlight CBDebug as a reliable and effective method for debugging based on real user feedback.
@@ 7 Acknowledgements
This research was supported by a gift to the LinkedIn–Cornell Bowers Strategic Partnership, and a grant from Infosys. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsors.
@@ A.1 Initial Training
  • MetaShift (Liang & Zou, 2022): A natural dataset where the spurious attribute is the scene context (indoor vs. outdoor). We use the version derived from the COCO dataset rather than Visual Genome, containing 2738 training samples.
@@ C.3 Augmentation Probabilities
We found $\gamma=2$ to work well in practice and leave extensive hyperparameter-tuning for future work.

### v2

@@ 1 Introduction
Concept Bottleneck Models (CBMs) (Koh et al., 2020) have emerged as a powerful architecture for interpretable vision classification. A CBM consists of two stages: a concept extractor first predicts a set of human-understandable concepts, which are then passed to an inference layer to produce the final label. This intermediate representation allows domain experts to inspect the model’s reasoning process and verify whether it aligns with their own. This capability is crucial in high-stakes domains, such as healthcare or scientific analysis, where errors are costly (Makary and Daniel, 2016) and expert validation is essential (Lekadir et al., 2025).
Beyond passive validation, CBMs enable test-time interventions (Koh et al., 2020; Espinosa Zarlenga et al., 2023). An expert can review the predicted concepts and directly correct them to influence the final prediction. For example, if a radiologist corrects a mispredicted concept, such as marking a lesion as present when the model missed it, the diagnosis may shift from benign to malignant. Such interventions elevate the expert from a passive auditor to an active participant in decision making. However, the effectiveness of these interventions hinges on a critical assumption: the learned concepts must align with expert knowledge. In practice, this alignment is often fragile. Data quality issues, sampling bias, and incomplete concept vocabularies can lead models to exploit spurious correlations or overlook key factors (Torralba and Efros, 2011; Geirhos et al., 2020). Because test-time edits address only surface errors, the same reasoning flaws inevitably reappear on new samples.
@@
In this work, we present a general debugging framework for CBMs that enables experts to globally edit a model’s reliance on undesired concepts, ensuring that its predictions are not only accurate but also right for the right reasons (Ross et al., 2017), and aligned with the domain expert’s reasoning. This framework follows a two-step process of Removal and Retraining (Figure 1).
@@ 2 Related Work
These models fall into two main architectural families. ProtoPNets (Chen et al., 2019) and their extensions (Nauta et al., 2023b; Ma et al., 2024; Carmichael et al., 2024), learn prototypical patches from the training set to make predictions. In this family, concept-level explanations take the form of representative image patches from the training data. VLM-CBMs (Debole et al., 2025) use a CLIP backbone to score the presence of textual concepts, as explored in many approaches (Yuksekgonul et al., 2023; Oikarinen et al., 2023; Srivastava et al., 2024), where concept-level explanations are the textual descriptions of each concept. Our framework can also extend to post-hoc XAI methods (e.g., interpreting neurons from class activation maps (Zhou et al., 2016) as concepts). In this work we focus on concept bottlenecks, which are interpretable-by-design, and evaluate one representative from each family: PIP-Net (Nauta et al., 2023b) and Post-hoc CBM (Yuksekgonul et al., 2023).
Interpretable Debugging. The goal of interpretable debugging is to enable a domain expert to interact with an interpretable model to detect and correct undesired behaviors. Early work focused on explanatory interactive learning (Ross et al., 2017; Teso and Kersting, 2019; Schramowski et al., 2020), with subsequent extensions to neuro-symbolic models (Stammer et al., 2022; 2021). Bontempelli et al. (2023) advanced these approaches, focusing on ProtoPNets and removing the need for a fixed concept vocabulary. Building on top of these works, we focus on a generalized unsupervised CBM architecture (Poeta et al., 2023).
Within unsupervised CBMs, spurious concept removal has been explored in both ProtoPNets and VLM-CBMs: Nauta et al. (2023a) study removal in PIP-Net on medical datasets, while Yuksekgonul et al. (2023); Rao et al. (2024); Chauhan et al. (2023) evaluate similar strategies for VLM-CBMs. Retraining efforts have focused mainly on ProtoPNets, such as adding object segmentation maps for extra supervision (Barnett et al., 2021) or ProtoPDebug (Bontempelli et al., 2023) which applies a forgetting loss on class-specific subsets of user-identified spurious concepts. Donnelly et al. (2025) targets task-specific spurious concepts and bypasses the need for retraining, but is limited to random patch selection for learning new concepts. For VLM-CBMs, Bontempelli et al. (2021) outline debugging strategies but lack empirical evaluation, while Hu et al. (2024) supports concept removal without the need for retraining through influence functions. We build on these directions with a general debugging framework for CBMs and an effective retraining approach, CBDebug, that leverages a novel connection between interpretable debugging and bias mitigation to remove undesired task-specific concepts and better align models with expert feedback.
@@
Supervised methods require auxiliary labels to reduce the impact of spurious correlations (Zheng and Makar, 2022; Sagawa et al., 2019; Makar et al., 2022; Kirichenko et al., 2022). While these approaches cannot be directly applied to unsupervised CBMs, CBDebug utilizes a crucial connection between interpretable debugging and these methods. By leveraging the model’s interpretability, we can bridge concept-level human feedback to sample-level auxiliary labels, effectively removing the reliance on any concept marked as undesired by the domain expert. This is accomplished by collecting the activation scores for each concept on every sample, which can then be used to form the auxiliary labels. Specifically, we instantiate our approach with permutation weighting (Arbour et al., 2021), which can handle high-dimensional, real-valued auxiliary labels. This allows us to use concept activations directly, unlike methods such as GroupDRO (Sagawa et al., 2019) that require discrete groups and would necessitate an additional clustering step to convert activation scores.
@@ 3.2 Removal
To guide concept removal, we adopt a simple binary feedback mechanism that assumes spuriosity is defined at the task level: each concept is either retained or marked for removal, based on expert input. This minimal supervision design ensures that our approach remains broadly applicable and easy to integrate into real-world workflows, where experts may have limited time or domain knowledge to provide detailed annotations, but does not directly account for class-specific spurious concepts.
@@ 4.2 Reweight
We adopt permutation weighting (Arbour et al., 2021), later applied to shortcut removal by Zheng and Makar (2022), to perform reweighting. Unlike group-based reweighting approaches such as GroupDRO (Sagawa et al., 2019), which require discrete group labels and often struggle with groups that have low support (necessitating clustering of $\smash{\mathbf{\widehat{V}}}$), permutation weighting naturally accommodates multi-dimensional, continuous auxiliary labels. By directly enforcing independence between $\smash{\mathbf{\widehat{V}}}$ and $Y$, it provides a more general and stable mechanism.
@@
where $\eta(y_{i},v_{i})$ denotes the estimated probability that a sample belongs to $D^{\prime}$. To ensure robust weights, we perform K-fold cross validation and average over multiple permutations. See Appendix C.1 for further analysis of effect of the number of folds and number of permutations on assigned sample weights per subgroup.
@@ 4.3 Augment
Samples assigned a low weight are more likely to be aligned with the bias we want to remove, and so we would like to focus our augmentation on these samples. To accomplish this, we can convert each sample weight $u_{i}$ into an augmentation probability $p_{aug}(x_{i})$. We first invert the sample weight $u_{i}$ by subtracting each weight from the maximum sample weight assigned. Then we normalize the resulting values to $[0,1]$ to convert them into probabilities, and raise the probabilities to a power $\gamma$ to increase contrast and reduce the likelihood of augmenting useful samples. See AppendixC.3 for further analysis on the impact of $\gamma$ on augmentation probabilities and downstream performance.
We then augment each sample with probability $p_{aug}(x_{i})$. Our augmentation strategy depends on the concept representation:
  • ProtoPNets: Since ProtoPNets operate on prototypical patches, we first randomly select $k$ spurious concepts from $\mathcal{C}_{spur}$ and perform CutMix (Yun et al., 2019) for each selected concept. To determine the actual patch used, we randomly choose from the top-ten highest-activated prototypical patches for that concept, following Nauta et al. (2023b)’s approach for visualizing concepts.
  • VLM-CBMs: Since VLM-CBMs use text-based concepts, we leverage a text-to-image-generated concept bank, following DISC (Wu et al., 2023). We randomly select a single spurious concept from $\mathcal{C}_{spur}$ and perform Mixup (Zhang, 2017) with an image of that concept randomly sampled from the concept bank.
This approach ensures that bias-aligned samples are more likely to be augmented, creating a richer training distribution that reduces spurious correlations without altering the true labels.
@@ 5 Experiments
Datasets. We use datasets with known spurious correlations: Waterbirds (Sagawa et al., 2019), MetaShift (Liang and Zou, 2022), CelebA (Liu et al., 2015), and ISIC (Codella et al., 2019). These datasets provide concrete testbeds for assessing how well CBDebug reduces reliance on undesired concepts, as their group structures allow performance to be measured directly across subpopulations.
@@ 5.1 Can CBDebug effectively retrain based on user feedback?
CBDebug. In contrast to these naive baselines, CBDebug improves worst-group accuracy by 7.5% on Waterbirds and 4.9% on MetaShift for PIP-Net, and by 26.1% on Waterbirds and 4.8% on MetaShift for Post-hoc CBM. CBDebug surpasses the previous state-of-the-art interpretable debugger, ProtoPDebug, while integrating its component steps into a framework that delivers more stable gains across settings.
These consistent improvements across architectures and datasets highlight CBDebug as a reliable and effective method for debugging based on real user feedback. We additionally compare CBDebug to unsupervised group-robustness methods in Appendix C.4, which rely on assumptions about how spurious correlations are learned, showing the benefit of human-guided debugging for performing bias mitigation.
@@ 6 Conclusions
7 Limitations
While CBDebug demonstrates consistent improvements across datasets and model architectures, several limitations remain. First, we observe substantial variance in performance metrics, particularly when using Post-hoc CBM as the backbone, where results are highly sensitive to initialization and training stochasticity. Although CBDebug reliably improves performance across settings, the effect sizes should be interpreted with this variability in mind.
Second, CBDebug’s effectiveness depends on the quality and consistency of concept feedback: inaccurate or adversarial feedback can undermine performance gains. Finally, our framework assumes task-level spurious concepts and therefore does not directly mitigate class-specific spurious correlations, which may limit applicability in settings where spuriosity varies across classes.
@@ 8 Acknowledgements
We thank Linxi Zhao and Sofian Zalouk for insightful discussions on the ideas and manuscript. This research was supported by a gift to the LinkedIn–Cornell Bowers Strategic Partnership, ARO grant W911NF-25-1-0254, BSF grant 2024101 and a grant from Infosys. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsors.
@@ A.1 Initial Training
  • MetaShift (Liang and Zou, 2022): A natural dataset where the spurious attribute is the scene context (indoor vs. outdoor). We use the version derived from the COCO dataset rather than Visual Genome, containing 2738 training samples.
@@ A.3 Retraining
A.4 Definitions
Permutation Weighting: A technique that reweights training samples to reduce reliance on a given auxiliary variable. Each sample receives a weight reflecting its likelihood under a permuted version of the auxiliary variable, encouraging the model to focus less on spurious correlations.
Worst-Group Accuracy (WGA): Measures the model’s performance on the group of data on which it performs the worst. This metric highlights whether the model fails on minority or atypical subgroups.
@@ C.3 Augmentation Probabilities
Table 8 shows worst-group accuracy (WGA) on Waterbirds for $\gamma\in\{1,2,5,10\}$. While there is some variability across seeds, WGA remains relatively stable, indicating that CBDebug is robust to the exact choice of $\gamma$. For Waterbirds, MetaShift, and CelebA, we use $\gamma=2$. For ISIC, we use $\gamma=5$ to reduce augmentation intensity and avoid distribution shifts caused by applying diffusion-generated concepts to medical images. Extensive tuning of this hyperparameter is left for future work.
Table 8: Worst-Group Accuracy across values of $\gamma$ (augmentation probability) for Post-hoc CBM on Waterbirds.
Average Worst
$\gamma=1$ 75.3 ±6.3 60.2 ±12.7
$\gamma=2$ 76.0 ±2.8 58.3 $_{\phantom{0}\pm 6.0}$
$\gamma=5$ 77.6 ±6.0 58.9 $_{\phantom{0}\pm 9.1}$
$\gamma=10$ 77.8 ±6.6 62.5 ±11.4